OpenStudy (xmoses1):

What does this mean: Evaluate C(7, 4) 35 210 5,040

OpenStudy (anonymous):

C(7,4) = C(7,3) because 3+4=7 =7*6*5 / (3*2*1) = 7*6*5 / 6 = 7*5 = 35

OpenStudy (xmoses1):

Oh...

OpenStudy (xmoses1):

That is actually so easy lol

OpenStudy (anonymous):

Im happy to help

OpenStudy (xmoses1):

Thanks @Mr.Question